OpenStudy (anonymous):

Find the avg rate of change of the function f(x)=x^2+5x from x1= 1 to x2=3

OpenStudy (anonymous):

rate of change is another way of saying slope. So use the 2 x-values given to determine the y values (f(x)). Then find the slope between those 2 points

OpenStudy (anonymous):

1/3?

OpenStudy (anonymous):

can you help me with one more CHmE

OpenStudy (anonymous):

you need to redo this problem. \[\text{slope}=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

what is the other y value?

OpenStudy (anonymous):

you find both by plugging in the 2 x-values given into the equation

OpenStudy (anonymous):

1 and 3

OpenStudy (anonymous):

\[y_1=(1)^2+5(1)\]\[y_2=(3)^2+5(3)\]

OpenStudy (anonymous):

y1=7 y2= 21

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

6, 24

OpenStudy (anonymous):

9

OpenStudy (anonymous):

correct !! Good job

OpenStudy (anonymous):

thanks!'